Lumi:

4.12 Unit 4 Test: Part 2

Lumi:

1. Melissa bought a new car with 12 optional features to the car. Each optional feature costs either $23.99 or $120.84. if she paid $675.28 for all the optional features, how many of each type did she buy? Write and solve an equation to solve this problem. Show all your work. x + y = 12 23.99x + 120.84y = 675.28 x = -y + 12 23.99(-y +12) + 120.84y = 675.28 -23.99y + 287.88 + 120.84y = 675.28 96.85y + 287.88 = 675.28 96.85y = 387.4 y = 4 x + 4 = 12 x = 8 Melissa bought 8 of the $23.99 features and 4 of the $120.84 features. 2. Ken mails three packages at the post office that weight 8.1, 8.5 and 8.9 pounds. The post office that Ken uses rounds the weight of a package to the nearest pound to determine the price of shipment. It costs $2 to mail a package that weighs 8 pounds and $2.75 for packages that weight 9 pounds. (a) How much did Ken pay to ship each package?(b) Ken has $7.25 in his wallet in cash. Use subtraction to determine whether Ken has enough to mail the three packages and how much extra money he has or needs. (a) After rounding, there is one "8 pound" package and two "9 pound" packages. To ship the first package, it costs $2 to ship the "8 pound" package and then a further $5.50 to ship the two "9 pound" packages. In total, it cost $7.50 to ship all the packages (b) 7.25 - 7.50 = -0.25 Ken does not have enough money to mail the three packages. He needs an additional $0.25 to mail the three packages.

Lumi:

If I am incorrect or you have any criticisms, please explain below.